OpenStudy (anonymous):

If f(x) = x2 and g(x) = x + 6, find g(f(0)).

OpenStudy (ivettef365):

first solve f(0), so substitute the x for 0 on f(x) function

OpenStudy (anonymous):

This is just saying put \( x^2 \) into into x + 6. So where ever there is an x in x + 6 substitute \( x^2 \) So solve for \( x^2 + 6 \)

OpenStudy (ivettef365):

not in this case that is when you are asked (gof)

OpenStudy (ivettef365):

so you have f(0) = 0^2 = 0 now substitute the 0 on the equation g(x)

OpenStudy (anonymous):

Yeah I miss read and did not see the 0 there

OpenStudy (ivettef365):

so you have 0 + 6 = 6